# SLA Management in Reconfigurable Multi-Agent RAG: A Systems Approach to Question Answering

### v1

@@ Abstract
Index Terms:
Multi-Agent Systems, Information Retrieval, Service Level Agreement, Quality of Service, Large Language Models, Retrieval Augmented Generation
I Introduction
Question Answering (QA) has been a pivotal area of research since the 1960s, aiming to provide precise answers to natural language questions. Unlike traditional search engines that return a list of hyperlinks to relevant pages, QA systems deliver direct answers to user queries [30].
The advent of generative Large Language Models (LLMs) has led search engines to integrate abstractive QA into their offerings—systems that generate answers which may not exist verbatim in source documents [4]. Additionally, new platforms specializing in abstractive QA, such as Perplexity [20] and ChatGPT Search [19], have emerged.
However, LLM-based abstractive QA exhibits notable limitations, including a lack of generalization to unseen data and a propensity for hallucination, i.e. the generation of false assertions [10]. Retrieval Augmented Generation (RAG) [8] addresses these shortcomings by combining LLMs with external knowledge sources, enabling the LLM to function as a reasoning layer rather than a static repository of facts. RAG retrieves context from external sources such as search engines, traditional databases, vector databases [11], or knowledge graphs [17, 28], and utilizes the LLM for reasoning and response generation.
@@ I Introduction
  • E-commerce Search: Highly sensitive to response times, where delays can significantly impact user experience and sales conversions [25, 1].
  • Customer Support: Requires not just correct answers but stylistically compliant ones that adhere to the brand-voice guidelines of the domain or company [18].
  • Legal and Healthcare Queries: Demand high answer quality due to the critical nature of the information, with less emphasis on cost or latency constraints [22].
@@
I-A Our Contributions
In this paper, we present a novel system-theoretic framework for SLA management in reconfigurable multi-agent RAG systems, specifically tailored for question-answering applications. Our main contributions are as follows:
@@ I-A Our Contributions
I-B Related Work
Scaling Retrieval-Augmented Generation (RAG) systems has been an active area of research, with most approaches focusing on vertical scaling—decomposing tasks into subtasks assigned to specialized components. Recent studies incorporate Large Language Models (LLMs) at the planning stages to mitigate issues like irrelevant retrieval and focus drift. For instance, PlanRAG uses an LLM to generate retrieval plans, enhancing context relevance [12]. Similarly, Retrieve-Plan-Generate (RPG) separates planning and answering stages to improve long-form question answering [14]. AutoGPT+P introduces planning loops for step-by-step plan generation in dynamic environments [2].
Horizontal scaling through multi-agent systems has also been explored, where multiple agents are synchronously assigned to the same question. LLM-Debate employs multiple LLM agents debating to reach better answers [5]. Chain-of-Thought Self-Consistency (CoT-SC) generates multiple reasoning paths and selects the most consistent answer [27]. Li et al. demonstrate that increasing the number of agents with simple voting improves performance [13].
Temporal scaling techniques involve sequential interactions with LLMs to enhance question-answering performance on complex queries. Shao et al. use iterative querying to refine answers for complex information needs [23]. Query refinement methods, such as RQ-RAG involve rewriting, decomposing, and disambiguating complex queries have also been shown to improve answer quality [3].
@@ I-B Related Work
Gao et al. introduce a modular RAG framework that encompasses horizontal, vertical, and temporal scaling techniques [9]. Zhang et al. propose an agentic information retrieval framework capable of handling complex queries by introducing IR agents with greater levels of autonomy [29]. However, these works lack experimental results on adapting systems to real-world environments.
Arbitration mechanisms have been proposed to improve decision-making in multi-agent systems. Some studies suggest using multiple LLMs to select the best response [26]. These methods improve answer quality but may not address trade-offs such as cost, latency, and security.
Evaluation frameworks like RAGAS assess RAG performance but may omit industry-relevant metrics such as adherence to stylistic constraints [6]. In practical applications, an answer must be not only correct but also stylistically appropriate.
@@
II Architecture
@@
II-A System Inputs and Outputs
User Query: The system accepts arbitrary input queries from users, such as: ”How do I reset my phone?”, ”How to reset phone”, or ”Burger joints near me”.
@@ II-A System Inputs and Outputs
Figure 1: A bird’s eye view of the architecture of the AbstractiveQA System
Task SLA/QoS Parameters: Specific requirements such as:
@@
Operational Environment Current conditions affecting the system, such as:
@@
The QA system’s outputs are Intent-Dependent Answers which consist of diverse responses tailored to the user’s intent and complying with specified SLAs.
II-B Intent Detection Module
The Intent Detection Module processes the user query to classify it into predefined intent categories, which may include:
@@ II-B Intent Detection Module
II-B1 Implementation Options
Implementation options for this module include various techniques, each with trade-offs in terms of computational cost, extensibility, and accuracy:
@@ II-C Planning Module
Figure 2: Possible Instantiation of Abstractive Question Answer Intent Handler
II-D Intent Handlers
@@
III Benchmark Dataset Creation and Metrics
III-A Dataset Creation
@@
  • Evaluation of the Intent Classifier (which is outside the scope of this paper).
@@
Each result was assigned with a total of three tags dependent on the availability of an answer in context, the provision of an answer, and the quality of the answer: Context Availability:
  • Answer Exists in Context: The agent’s retrieval process returned at least one result from which a correct direct answer could be sourced.
  • Answer Does Not Exist in Context: The agent’s retrieval process did not return any results from which a correct direct answer could be sourced.
Answer Provision:
  • Answer Provided: An answer was returned.
  • No Answer Provided: The returned answer was null.
Answer Quality:
  • Correct Answer: The returned answer was correct and aligned with the task’s non-functional requirements (e.g., adhered to the specified style in the prompt).
  • Hallucination: The returned answer was not grounded in the provided context.
  • Incongruent Response: The returned answer did not align with the task’s non-functional requirements.
III-B Metrics
III-B1 Answer Quality Metrics
@@
  • Recall: The number of correct answers divided by the number of times the answer exists in the global context (the union of all contexts from all agents). [^1] The use of the global context to compute precision and recall introduces limitations to our dataset, as the exact upper bound on the number of answers in any context is unknown and could not be determined within our time constraints. However, we found that the difference in the number of answers sourced from the global context between the four-agent and five-agent systems was minimal (one to two instances, depending on configuration). This suggests that the overall impact is likely negligible with the retrieval mechanisms used.
  • F1 Score: The harmonic mean of precision and recall.
@@
  • Hallucination Rate: The number of hallucinations divided by the total number of answers provided, where a hallucination is any answer not grounded in the context.  [^2] The Hallucination Rate is analogous to the False Positive Rate, with the key difference being that false positives in this context include answers not grounded in the provided context (extrinsic hallucinations), contradicted by the provided context (intrinsic hallucinations), or incongruent responses. Further details on categories of hallucination can be found in [10].
@@
Precision, Recall, and F1: In traditional information retrieval, precision is defined as the proportion of retrieved documents that are relevant, and recall is the proportion of relevant documents that are retrieved [16]. In our question answering domain, we adapted these definitions to suit our specific context.
Incongruent Responses: For many users of this system, particularly brands, it is insufficient to simply return a correct answer; the answer must also adhere to the non-functional and stylistic requirements specified in the prompt. This includes guidelines such as not breaking the fourth wall or referencing the context directly. Ensuring compliance with these requirements is essential for maintaining brand voice and user experience.
III-B2 System Cost
The expected cost of our QA system, as used in our experiments, is calculated as the sum of the fixed system overhead and the costs associated with QA agents and arbitration, aggregated across all intents. The formula is:
$C_{\text{sys}}=C_{\text{overhead}}+\sum_{\text{intent}=1}^{I}\big{(}C_{\text{agent}}\cdot N_{\text{intent}}+C_{\text{arbitration}}(N_{\text{intent}})\big{)}$
@@ III-B2 System Cost
  • $C_{\text{sys}}$: Total expected system cost.
  • $C_{\text{overhead}}$: Fixed system overhead cost.
  • $C_{\text{agent}}$: Cost per QA agent.
  • $N_{\text{intent}}$: Number of agents for a specific intent.
  • $C_{\text{arbitration}}(N_{\text{intent}})$: Arbitration cost for $N_{\text{intent}}$ agents.
  • $I$: Total number of intents.
This equation models the specific cost structure of our QA system, where both the agent and arbitration costs are functions of the number of agents configured for each intent.
III-B3 System Latency
The system latency for our implementation is defined as the sum of the fixed overhead latency, the maximum latency of the QA agents configured for a given intent (since agent operations are asynchronous), and the arbitration latency for those agents. Mathematically, this is expressed as:
$L_{\text{sys}}=L_{\text{overhead}}+\max_{i\in\text{agents}}(L_{\text{QA},i})+L_{\text{arbitration}}(N_{\text{intent}})$
@@ III-B3 System Latency
  • $\max_{i\in\text{agents}}(L_{\text{QA},i})$: Maximum latency among all QA agents included in the intent service.
  • $L_{\text{arbitration}}(N_{\text{intent}})$: Arbitration latency for $N_{\text{intent}}$ agents.
This formula reflects the asynchronous nature of the QA agent operations and the specific latency structure of our experimental implementation.
IV Agent Implementation
@@
Figure 3: The RAG QA Agent Architecture we use in our experiments
IV-A Agent Architecture
The architecture of an individual QA agent used in our experiments, shown in Figure 3, is designed to be flexible and configurable based on the output of the Planning Module. The specific parameterization of each agent depends on factors such as which backend data sources are provisioned and caching strategies. While the implementation details of the Planning Module are outside the scope of this paper, we provide an overview of the QA agent’s architecture and workflow.
@@ IV-A Agent Architecture
  • Reasoning Module: Powered by GPT-4 [7]. Generates answers based on the user query and preprocessed context. Details of the prompt used are discussed in Section IV-C.
Figure 4: Architecture of the Vector Search Process we use in our experiments
IV-B Retrieval Module Details
@@
  • Embedding Model: Documents are embedded using a transformer model based on MPNet [24]. Embeddings capture semantic information for effective similarity comparisons.
  • Vector Database: Indexed using Hierarchical Navigable Small World (HNSW) graphs [15] for efficient approximate nearest neighbor search. At query time, the same embedding model is used to encode the user query. The resulting query vector $\mathbf{q}$ is used to retrieve documents whose vectors $\mathbf{d}$ are similar, based on cosine similarity:
@@
Prompt Example prompt = f""" Context: {context_entities}. Use only the provided context to answer the question: "{query}" to the best of your ability and in a few sentences or less. If there is more than one answer, summarize the options. Provide the "uid" values of the object you used to inform the answer (do NOT use the "id" value) Return the answer in a json dictionary format [Example: {"answer":"This is the answer", "uid_list": [12345, 98342]}] Do not refer to the context in your answer. """
Figure 5: An example of the constructed prompt used by the QA Agent’s Reasoning Module.
IV-C Prompt Engineering
After re-ranking and/or thresholding, we construct the prompt for the LLM. The preprocessed list of context documents and the original user query are interpolated into the prompt at appropriate places. An example prompt is shown in Figure 5:
V Experiments
We conducted experiments on individual QA agents by varying the post-retrieval preprocessing logic used to prepare the input prompts for the GPT-4-powered Reasoning Module. Given the LLM’s context window limitation of 8,000 tokens, we needed to prune the retrieved context appropriately. We developed several pruning methods aiming to balance content relevance with the token limit. For each agent, we measured Answer Quality using the metrics described in Section III-B1.

### v2

@@ Abstract.
Retrieval Augmented Generation, Multi-Agent Systems, Service Level Agreement
ccs: Information systems Information retrievalccs: Applied computing Service-oriented architecturesccs: Computing methodologies Distributed artificial intelligence
1. Introduction
Question Answering (QA) has been a pivotal area of research since the 1960s, aiming to provide precise answers to natural language questions. Unlike traditional search engines that return a list of hyperlinks to relevant pages, QA systems deliver direct answers to user queries (Zhu et al., 2021).
The advent of generative Large Language Models (LLMs) has led search engines to integrate abstractive QA into their offerings—systems that generate answers which may not exist verbatim in source documents (Danny D’Cruze, 2024). Additionally, new platforms specializing in abstractive QA, such as Perplexity (Perplexity AI, 2024) and ChatGPT Search (OpenAI, 2024), have emerged.
However, LLM-based abstractive QA exhibits notable limitations, including a lack of generalization to unseen data and a propensity for hallucination, i.e. the generation of false assertions (Ji et al., 2023). Retrieval Augmented Generation (RAG) (Gao et al., 2024a) addresses these shortcomings by combining LLMs with external knowledge sources, enabling the LLM to function as a reasoning layer rather than a static repository of facts. RAG retrieves context from external sources such as search engines, traditional databases, vector databases (Jing et al., 2024), or knowledge graphs (Martino et al., 2023; Xu et al., 2024), and utilizes the LLM for reasoning and response generation.
@@ 1. Introduction
  • E-commerce Search: Highly sensitive to response times, where delays can significantly impact user experience and sales conversions (van Vessum, nd; Achary, 2023).
  • Customer Support: Requires not just correct answers but stylistically compliant ones that adhere to the brand-voice guidelines of the domain or company (Murár et al., 2024).
  • Legal and Healthcare Queries: Demand high answer quality due to the critical nature of the information, with less emphasis on cost or latency constraints (Samat, 2024).
Furthermore, because modern cloud providers (e.g., SaaS and PaaS platforms) host these varied workloads side by side on shared infrastructure, the system’s overall utility is increased: by optimally provisioning and multiplexing resources across collocated services with diverse SLA requirements, we achieve higher utilization and more efficient service delivery.
@@
1.1. Our Contributions
We present a novel system-theoretic framework for SLA management in reconfigurable multi-agent RAG systems, specifically tailored for question-answering applications. Our main contributions are as follows:
@@ 1.1. Our Contributions
1.2. Related Work
Scaling Retrieval-Augmented Generation (RAG) systems has been an active area of research, with most approaches focusing on vertical scaling—decomposing tasks into subtasks assigned to specialized components. Recent studies incorporate Large Language Models (LLMs) at the planning stages to mitigate issues like irrelevant retrieval and focus drift. For instance, PlanRAG uses an LLM to generate retrieval plans, enhancing context relevance (Lee et al., 2024). Similarly, Retrieve-Plan-Generate (RPG) separates planning and answering stages to improve long-form question answering (Lyu et al., 2024). AutoGPT+P introduces planning loops for step-by-step plan generation in dynamic environments (Birr et al., 2024).
Horizontal scaling through multi-agent systems has also been explored, where multiple agents are synchronously assigned to the same question. LLM-Debate employs multiple LLM agents debating to reach better answers (Du et al., 2023). Chain-of-Thought Self-Consistency (CoT-SC) generates multiple reasoning paths and selects the most consistent answer (Wang et al., 2023). Li et al. demonstrate that increasing the number of agents with simple voting improves performance (Li et al., 2024).
Temporal scaling techniques involve sequential interactions with LLMs to enhance question-answering performance on complex queries. Shao et al. use iterative querying to refine answers for complex information needs (Shao et al., 2023). Query refinement methods, such as RQ-RAG involve rewriting, decomposing, and disambiguating complex queries have also been shown to improve answer quality (Chan et al., 2024).
@@ 1.2. Related Work
Gao et al. introduce a modular RAG framework that encompasses horizontal, vertical, and temporal scaling techniques (Gao et al., 2024b). Zhang et al. propose an agentic information retrieval framework capable of handling complex queries by introducing IR agents with greater levels of autonomy (Zhang et al., 2024). However, these works lack experimental results on adapting systems to real-world environments.
Arbitration mechanisms have been proposed to improve decision-making in multi-agent systems. Some studies suggest using multiple LLMs to select the best response (Verga et al., 2024). These methods improve answer quality but may not address trade-offs such as cost, latency, and security.
Evaluation frameworks like RAGAS assess RAG performance but may omit industry-relevant metrics such as adherence to stylistic constraints (Es et al., 2023). In practical applications, an answer must be not only correct but also stylistically appropriate.
@@
2. SLA Composition in Multi-Agent QA Systems
We adopt a compositional perspective on Service Level Agreements (SLAs), decomposing a complex SLA into multiple interacting Quality-of-Service (QoS) attributes—such as response latency, factual accuracy, stylistic adherence, and cost—and then combining them into an overall service commitment. Under this framework, the composite SLA for a QA system is defined as
$\mathrm{SLA}_{\mathrm{composite}}=f(q_{1},q_{2},\dots,q_{k}),$
where each $q_{i}$ denotes a QoS attribute (e.g., latency, answer quality, stylistic compliance) and the function $f$ captures how these attributes interact.
In our reconfigurable multi-agent RAG architecture, every module—from Intent Detection and Planning to individual QA agents and the Arbitration component—contributes toward one or more of these QoS attributes. For example, the Planning Module dynamically allocates both the number and configuration of RAG agents based on the incoming query’s requirements and current operational conditions; this allocation explicitly considers the relevant SLA sub-attributes to ensure the final answer meets its targets for speed, accuracy, and style.
A continuous monitoring subsystem then audits each module’s compliance with its assigned sub-SLAs in real time. Whenever a deviation from the target QoS levels is detected—whether in response time, factual correctness, or stylistic conformity—adaptive reconfiguration is triggered. Such adjustments might include tuning the arbitration thresholds, shifting workload between cloud and local LLM instances, or reallocating compute resources. This feedback loop ensures that the composite SLA is upheld by continuously reconciling the performance of individual components.
By embedding this compositional SLA framework into our multi-agent QA system, we achieve a systematic balance among answer quality, latency, and cost—yielding a robust, adaptable service that satisfies diverse and stringent SLA requirements in real-world deployments.
3. Architecture
@@
3.1. System Inputs and Outputs
3.1.1. User Query
: The system accepts arbitrary input queries from users, such as: “How do I reset my phone?”, “How to reset phone”, or “Burger joints near me”.
@@ 3.1.1. User Query
Figure 1. A bird’s eye view of the architecture of the AbstractiveQA System
3.1.2. Task SLA/QoS Parameters
: Specific requirements such as:
@@ 3.1.2. Task SLA/QoS Parameters
3.1.3. Operational Environment
Current conditions affecting the system, such as:
@@ 3.1.3. Operational Environment
3.1.4. Results
: The QA system’s outputs are Intent-Dependent Answers which consist of diverse responses tailored to the user’s intent and complying with specified SLAs.
3.2. Intent Detection Module
The Intent Detection Module processes the user query to classify it into predefined intent categories, focusing on determining the what of the input query [^1] The what refers to the type or kind of query being made, as opposed to the why, which pertains to the motivation behind the user’s question., i.e., what kind of query it is:
@@ 3.2. Intent Detection Module
3.2.1. Implementation Options
Implementation options for this module include various techniques, each with trade-offs in terms of computational cost, extensibility, and accuracy [^2] In this context, accuracy is used in a qualitative sense to measure the quality of outcomes, rather than the quantitative definition commonly used in machine learning.:
@@ 3.3. Planning Module
Figure 2. Possible Instantiation of Abstractive Question Answer Intent Handler
3.4. Intent Handlers
@@
4. Benchmark Dataset Creation and Metrics
4.1. Dataset Creation
@@
  • Evaluation of the Intent Classifier (which is outside the scope of this research).
@@
Each result was assigned three attributes reflecting: (i) the availability of an answer in context, (ii) the provision of an answer, and (iii) the quality of the answer. Detailed definitions for these attributes are provided in Table 1.
Category Option Description
Context Availability Answer Exists in Context At least one agent’s retrieval process returned a result from which a correct direct answer could be derived.
Answer Does Not Exist in Context None of the agents’ retrieval processes returned a result from which a correct direct answer could be derived.
Answer Provision Answer Provided An answer was returned (regardless of its correctness).
No Answer Provided A negative, or null, answer was returned.
Answer Quality Correct Answer The returned answer was correct and aligned with the task’s non-functional requirements (e.g., adhered to the specified style in the prompt).
Hallucination The returned answer was not grounded in the provided context.
Incongruent Response The returned answer did not adhere to the task’s non-functional requirements.
Table 1. Summary of Answer Annotation Attributes
4.2. Metrics
4.2.1. Answer Quality Metrics
@@
  • Recall: The number of correct answers divided by the number of times the answer exists in the global context (i.e., the union of all contexts from all agents). [^3] The use of the global context to compute precision and recall introduces limitations to our dataset, as the exact upper bound on the number of answers in any context is unknown and could not be determined within our time constraints. However, we found that the difference in the number of answers sourced from the global context between the four-agent and five-agent systems was minimal (one to two instances, depending on configuration). This suggests that the overall impact is likely negligible with the retrieval mechanisms used.
  • F1 Score: The harmonic mean of precision and recall:
@@
  • Hallucination Rate: The number of hallucinations divided by the total number of answers provided, where a hallucination is any answer not grounded in the context. [^4] The Hallucination Rate is analogous to the False Positive Rate, with the key difference being that false positives in this context include answers not grounded in the provided context (extrinsic hallucinations), contradicted by the provided context (intrinsic hallucinations), or incongruent responses. Further details on categories of hallucination can be found in (Ji et al., 2023).
@@
Precision, Recall, and F1:
In traditional information retrieval, precision is defined as the proportion of retrieved documents that are relevant, and recall as the proportion of relevant documents that are retrieved (Manning et al., 2008). In our question answering domain, we adapted these definitions to suit our specific context.
Incongruent Responses:
For many users of this system, particularly brands, it is insufficient to simply return a correct answer; the answer must also adhere to the non-functional and stylistic requirements specified in the prompt. This includes guidelines such as not breaking the fourth wall or referencing the context directly. Ensuring compliance with these requirements is essential for maintaining brand voice and user experience.
4.2.2. System Cost
Our model incorporates a fixed overhead cost per question, the cumulative cost of the diverse QA agents deployed for the intent, and the associated arbitration cost. Mathematically, this is expressed as:
$C_{\text{sys}}=C_{\text{overhead}}+\sum_{i=1}^{N_{\text{intent}}}C_{\text{agent},i}+C_{\text{arbitration}}(N_{\text{intent}})$
@@ 4.2.2. System Cost
  • $C_{\text{sys}}$: Total system cost of answering one question.
  • $C_{\text{overhead}}$: Fixed system overhead cost per question.
  • $C_{\text{agent},i}$: Cost of the $i$-th QA agent.
  • $N_{\text{intent}}$: Number of QA agents deployed for the intent.
  • $C_{\text{arbitration}}(N_{\text{intent}})$: Arbitration cost for the $N_{\text{intent}}$ agents.
4.2.3. System Latency
The system latency for our implementation is defined as the sum of the fixed overhead latency, the maximum latency among the QA agents configured for a given intent (since agent operations are asynchronous and parallel), and the arbitration latency for those agents. Mathematically, this is expressed as:
$L_{\text{sys}}=L_{\text{overhead}}+\max_{1\leq i\leq N_{\text{intent}}}L_{\text{QA},i}+L_{\text{arbitration}}(N_{\text{intent}})$
@@ 4.2.3. System Latency
  • $\max_{1\leq i\leq N_{\text{intent}}}L_{\text{QA},i}$: Maximum latency among all $N_{\text{intent}}$ QA agents.
  • $L_{\text{arbitration}}(N_{\text{intent}})$: Arbitration latency for the $N_{\text{intent}}$ agents.
5. Agent Implementation
@@
Figure 3. The RAG QA Agent Architecture we use in our experiments
5.1. Agent Architecture
The architecture of an individual QA agent used in our experiments, shown in Figure 3, is designed to be flexible and configurable based on the output of the Planning Module. The specific parameterization of each agent depends on factors such as which backend data sources are provisioned and caching strategies. While the implementation details of the Planning Module are outside the scope of this research, we provide an overview of the QA agent’s architecture and workflow.
@@ 5.1. Agent Architecture
  • Reasoning Module: Powered by GPT-4 (et al., 2024). Generates answers based on the user query and preprocessed context. Details of the prompt used are discussed in Section 5.3.
Figure 4. Architecture of the Vector Search Process we use in our experiments
5.2. Retrieval Module Details
@@
  • Embedding Model: Documents are embedded using a transformer model based on MPNet (Song et al., 2020). Embeddings capture semantic information for effective similarity comparisons.
  • Vector Database: Indexed using Hierarchical Navigable Small World (HNSW) graphs (Malkov and Yashunin, 2016) for efficient approximate nearest neighbor search. At query time, the same embedding model is used to encode the user query. The resulting query vector $\mathbf{q}$ is used to retrieve documents whose vectors $\mathbf{d}$ are similar, based on cosine similarity:
@@
5.3. Prompt Engineering
After re-ranking and/or thresholding, we construct the prompt for the LLM. The preprocessed list of context documents and the original user query are interpolated into the prompt at appropriate places. An example prompt is shown in Figure 5.
6. Experiments
We conducted experiments on individual QA agents by varying the post-retrieval preprocessing logic used to prepare the input prompts for the GPT-4-powered Reasoning Module. Given the LLM’s context window limitation of 8,000 tokens, we needed to prune the retrieved context appropriately. We developed several pruning methods aiming to balance content relevance with the token limit. For each agent, we measured Answer Quality using the metrics described in Section 4.2.1.
Prompt Example Context: {context_entities}. Use only the provided context to answer the question: ‘{query}’ to the best of your ability and in a few sentences or less. If there is more than one answer, summarize the options. Provide the uid values of the object you used to inform the answer (do NOT use the id value). Return the answer in a json dictionary format [Example: { answer’: This is the answer’, uid_list’: [12345, 98342] }] Do not refer to the context in your answer.
Figure 5. An example of the constructed prompt used by the QA Agent’s Reasoning Module.